Natalia spent $28.10 at an office supply store when she purchased boxes of pens and paper clips. The pens cost $3.50 per box, and the paper clips cost $0.70 per box. How many boxes of each did she buy if she purchased 10 items all together?
Answer by richwmiller(17219)

There is something wrong with the problem
a+b=10,
3.5*a+0.7*b=28.1
a=10-b
3.5*(10-b)+0.7*b=28.1
35.0-3.5b+0.7*b=28.1
-2.8*b=-6.9
b=2.46428571
a=10-b
a=7.53571429 at $3.5 and b=2.46428571 at $0.7
check
7.53571429+2.46428571=10
3.5*a+0.7*b=28.1
3.5*7.53571429+0.7*2.46428571=28.1
26.375+1.725=28.1
28.1=28.1
ok